An electron enters a region of space containing a uniform 0.0000281-T magnetic field. Its speed is 121 m/s and it enters perpendicularly to the field. Under these conditions, the electron undergoes circular motion Find the radius r of the electrons path, and the frequency fof the motion. Number r- Number Hz f-
Answer #1

The force experienced by the particle when it enters perpendicular to the field F = qvB

where q is the charge of the particle( here, charge of electron), v is the velocity of the particle and B is magnetic field or magnetic flux density.

F-1.6 * 10-19C * 121m/s * 0.000028 1T= 544.016 * 10-24M

When in circular motion, acceleration = v^{2}/R where v is the velocity and R is the radius of circular motion.

We know that F = ma where m is the mass of particle and a is the accelartion(mass of electron is 9.1 * 10-31 kg).

Thus, F = m * v^{2}/R = 544.016 * 10^{-24} N

R = m * v2/F = 24.4 * 10-18 m = 24.4 attometer

The frequency of particle in circular motion due to magnetic field is called cyclotron freuqency.

This frequency = f = qB/ (2\pi m) = 1.6 * 10^{-19}*0.0000281/ (2\pi * 9.1 * 10^{-31})

= 786330.35 Hz